# State-action control barrier functions: Imposing safety on learning-based control with low online computational costs

### v1

@@ Abstract
Learning-based control with safety guarantees usually requires real-time safety certification and modifications of possibly unsafe learning-based policies. The control barrier function (CBF) method uses a safety filter containing a constrained optimization problem to produce safe policies. However, finding a valid CBF for a general nonlinear system requires a complex function parameterization, which in general, makes the policy optimization problem difficult to solve in real time. For nonlinear systems with nonlinear state constraints, this paper proposes the novel concept of state-action CBFs, which not only characterize the safety at each state but also evaluate the control inputs taken at each state. State-action CBFs, in contrast to CBFs, enable a flexible parameterization, resulting in a safety filter that involves a convex quadratic optimization problem. This, in turn, significantly alleviates the online computational burden. To synthesize state-action CBFs, we propose a learning-based approach exploiting Hamilton-Jacobi reachability. The effect of learning errors on the effectiveness of state-action CBFs is addressed by constraint tightening and introducing a new concept called contractive CBFs. These contributions ensure formal safety guarantees for learned CBFs and control policies, enhancing the applicability of learning-based control in real-time scenarios. Simulation results on an inverted pendulum with elastic walls validate the proposed CBFs in terms of constraint satisfaction and CPU time.
@@ I Introduction
Learning-based control methods have demonstrated extensive success across many applications, such as autonomous vehicles [1] and robot manipulation [2]. Both supervised learning and reinforcement learning (RL) have been widely used for controller synthesis. However, learning-based controllers may provide unsafe control actions that result in undesirable or even destructive effects on the system. In control systems, safety means that the trajectories of the closed-loop system should satisfy state and input constraints for the entirety of the system’s evolution. The lack of safety guarantees limits the ability of learning-based control to achieve safety-critical tasks.
There has been an increasing interest in designing strategies to make learning-based controllers safe. Among the available methodologies including penalty methods [3], constraint elimination [4], and using invariant sets [5, 6], using control-barrier functions (CBFs) [7, 8] for safe controller synthesis is a universal and convenient approach. CBFs are energy-based functions and use their sub-level sets to characterize the safety of the dynamical system, avoiding the use of complex invariant set representations. Given any learning-based controller without safety considerations, an online optimization problem, called a safety filter, is then solved to find the closest safe input to this controller, while safety is gained by adding a constraint based on a CBF into the optimization problem.
@@
In the realm of CBF-based methodologies, a prominent limitation is the absence of universally applicable techniques for generating valid CBFs, thereby necessitating reliance on manually designed or problem-specific CBFs. For some particular kinds of systems such as linear, piecewise affine [9], or polynomial systems [10], CBFs can be computed by solving convex optimization problems. For nonlinear systems with general constraints, using learning-based algorithms accompanied by advanced function approximators to estimate CBFs has been explored in several contexts. Roughly, there are four kinds of methods to learn CBFs: the optimization-based method [11], the learner-verifier method [12], the Hamilton-Jacobi (HJ) reachability method [13], and the predictive safety filter method [14, 15]. The first two methods can be conservative, i.e., the learned safe set can be a small subset of the maximal controlled invariant set. In comparison, the HJ reachability method relies on the dynamic programming principle to approximate the maximal controlled invariant set iteratively, while the predictive safety filter method is based on a receding-horizon open-loop optimal control problem to implicitly determine the safe set, which also converges to the maximal controlled invariant set as the horizon goes to infinity [16]. A comprehensive comparison of HJ reachability, predictive safety filters, and CBFs can be found in [17].
@@
No matter how the CBF is learned, to control a given system with safety guarantees, it is inevitable to solve an online optimization problem with CBF-based constraints, which is usually non-convex for discrete-time nonlinear systems [18]. To obtain a satisfactory approximation accuracy, sophisticated function approximators such as deep neural networks (NNs) are commonly used to parameterize the CBF [14, 7]. However, this will inherently cause non-convexity and increase the complexity of the optimization problem. As a result, the increased online computational load makes the CBF-based approach unsuitable for situations where fast computation of control inputs is required. Besides, the effect of approximation errors on the validity of the learned CBFs and the resulting control policies has not yet been addressed, as highlighted in [7].
This paper contributes to the state of the art in the following aspects. To impose safety on nonlinear systems at low online computational costs, we propose the novel concept of state-action CBFs, the zero sub-level sets of which can be used in the safety filter to generate a safe policy. Compared to standard CBFs, state-action CBFs allow for a flexible parameterization so that the resulting safety filter contains a convex quadratic optimization problem. Consequently, the online computation burden is greatly eased. To analyze the effect of approximation errors on learned CBFs, we propose a constraint tightening approach and introduce a novel concept of contractive CBFs. When learning such CBFs, the invariance property of CBFs can be retained, despite the presence of approximation errors. We also explore the connection between state-action CBFs and (contractive) CBFs, and develop a new learning-based method to approximate state-action CBFs. As a result, the safety of the policy filtered by state action CBFs is preserved for a sufficiently small approximation error.
@@ II-A Preliminaries
where $x_{t}\in\mathcal{X}\subseteq\mathbb{R}^{n_{x}}$ and $u_{t}\in\mathcal{U}\subseteq\mathbb{R}^{n_{u}}$ are the state and the input at time step $t$, and $f(\cdot,\cdot):\mathcal{X}\times\mathcal{U}\to\mathcal{X}$ is a continuous function satisfying $f(0,0)=0$. We consider a constrained optimal control problem in which the states and inputs should satisfy time-invariant constraints: $x_{t}\in X:=\{x\in\mathcal{X}|h(x)\leq 0\}$ and $u_{t}\in U\subseteq\mathcal{U}$. Here, $h(\cdot):\mathcal{X}\to\mathbb{R}$ is a continuous function that defines the state constraint [^1] For the constraint defined by multiple inequalities $h_{i}(x)\leq 0,\;i=1,2,...,I$, we can let $h(x)=\max_{i\in\{1,...,I\}}h_{i}(x)$. The set $\{x|h(x)\leq 0\}$ is then identical to $\{x|h_{i}(x)\leq 0,\;i=1,2,...,I\}$, and $h$ will be continuous if each $h_{i}$ is continuous.. Throughout the paper, we assume that we completely know $f$. Besides, we assume that $X$ is compact and that $U$ is a polytope.
The control objective is to regulate a redefined control policy, e.g., a policy learned by reinforcement learning or a policy that approximates model predictive control (MPC) [19], so that the safety of the system is ensured.
@@ II-B Control barrier functions
$\min_{u\in U}B(f(x,u))\leq 0,\;\forall x\in\mathcal{S}_{B}.$ (3)
Furthermore, a CBF is called an exponential CBF if it satisfies (2) with 0 replaced by $B(x)$ and if there exists a $\beta\in(0,1)$ such that
$\min_{u\in U}B(f(x,u))\leq\beta B(x),\;\forall x\in\mathcal{S}_{B}.$ (4)
Condition (2) is equivalent to $\mathcal{S}_{B}\subseteq X$, which is usually assumed in literature [18]. With a CBF available, one generate a safe control policy in $\mathcal{S}_{B}$ by using the following optimization-based approach:
$\pi_{\mathrm{safe}}(x)=\arg\min_{u\in U}$ $\;||u-\pi_{0}(x)||_{2}$
@@
[…]
@@
For general nonlinear systems with state and input constraints, synthesizing a non-conservative CBF is a difficult task. To deal with this issue, using advanced function approximators to learn a CBF certificate has received much attention. One can see [7] for a comprehensive survey.
II-C Problem formulation
Problem P1: high online computational complexity. One main limitation of (II-B) is that it needs to solve a usually non-convex optimization problem in real time. If a CBF is represented by a complex function approximator $B_{\theta}(\cdot)$ with the parameter $\theta$, solving (II-B) may take much online computation time and result in very sub-optimal solutions. Even if a convex $B_{\theta}$ is formed, the constraint in (II-B) could be non-convex due to the nonlinearity of $f$.
Problem P2: effects of approximation errors. Another limitation of (II-B) is that the approximation error of $B(\cdot)$ may affect the safety of the system controlled by the optimizer of (II-B) with $B(\cdot)$ replaced by $B_{\theta}(\cdot)$. Besides, the recursive feasibility of (II-B) is also not guaranteed with $B_{\theta}(\cdot)$.
@@ III State-action control barrier function
where $Q(\cdot,\cdot):\mathcal{X}\times U\to\mathbb{R}$ is a function of states and actions. We will analyze how to enforce the safety of $\pi_{\mathrm{safe}}$ by imposing conditions on $Q$. To achieve this, we introduce the definition of state-action CBFs as follows.
@@
A continuous function $Q(\cdot,\cdot):\mathcal{X}\times U\to\mathbb{R}$ is called a state-action control barrier function with a corresponding safe set $\mathcal{S}_{Q}:=\{x\in X|\min_{u\in U}Q(x,u)\leq 0\}$, if $\mathcal{S}_{Q}$ is non-empty, and for any $x\in\mathcal{S}_{Q}$, any $u\in U$ that satisfies $Q(x,u)\leq 0$ will make $f(x,u)\in\mathcal{S}_{Q}$.
With the definition of state-action CBF, we have the following properties:
@@
(i) Consider the safety filter (6), any state-action CBF $Q$ will render $\pi_{\mathrm{safe}}$ safe in $\mathcal{S}_{Q}$. In other words, $\mathcal{S}_{Q}$ is control-invariant. (ii) If $B$ is a CBF, $Q(\cdot,\cdot)=B(f(\cdot,\cdot))$ will be a state-action CBF and the safe sets satisfy $\mathcal{S}_{B}\subseteq\mathcal{S}_{Q}$.
@@
Similarly to designing CBF, the challenge of using (6) is that the explicit form of a state-action CBF cannot be directly obtained based on Definition 3. However, as we know that $Q(\cdot,\cdot)=B(f(\cdot,\cdot))$ is a state-action CBF if $B$ is a CBF, if the exact value $B(x^{\mathrm{s}})$ of a CBF on any state sample $x^{\mathrm{s}}$ is available, we can use supervised learning-based methods to approximate $Q$, using a parameterized function approximator $Q_{\theta}$ with the parameter $\theta$.
@@
which will make problem (6) a convex QP with linear and convex quadratic constraints. In (7), $q_{1,\theta}(\cdot):\mathbb{R}^{n_{x}}\to\mathbb{R}$, $q_{2,\theta}(\cdot):\mathbb{R}^{n_{x}}\to\mathbb{R}^{n_{x}}$, and $Q_{3,\theta}(\cdot):\mathbb{R}^{n_{x}}\to\mathbb{R}^{n_{x}\times n_{x}}$ are parameterized functions with all parameters condensed in $\theta$.
IV Constraint tightening and contractive CBF
To cope with Problem P2, we propose a method to compute conservative CBFs based on state constraint tightening. In particular, we consider two kinds of conservative CBFs. The first kind includes the CBFs of system (1) under the tightened state constraint $X_{\lambda}:=\{x\in\mathcal{X}|h(x)+\lambda\leq 0\}$ where $\lambda$ is a positive constant. In this situation, the CBFs still satisfy (3) or (4), while condition (2) is tightened to $h(x)+\lambda\leq B(x),\forall x\in\mathcal{S}_{B}$.
The second one, called the contractive CBF, is defined as follows
Definition 4 (Contractive CBF)
A CBF $B(\cdot):\mathcal{X}\to\mathbb{R}$ is called a $\lambda$-contractive CBF [^2] Note that in fact the definition of contractive CBFs given here is not consistent with the common definition of contraction mappings [20]. Formally speaking, the CBF we define should be called a CBF with a contractive safe set [21]. However, for compactness we adopt “contractive CBF” in the paper. if there exists a $\lambda>0$ such that

### v2

@@ Abstract
Learning-based control with safety guarantees usually requires real-time safety certification and modifications of possibly unsafe learning-based policies. The control barrier function (CBF) method uses a safety filter containing a constrained optimization problem to produce safe policies. However, finding a valid CBF for a general nonlinear system requires a complex function parameterization, which in general makes the policy optimization problem difficult to solve in real time. For nonlinear systems with nonlinear state constraints, this paper proposes the novel concept of state-action CBFs (SACBFs), which do not only characterize the safety at each state but also evaluate the control inputs taken at each state. SACBFs, in contrast to CBFs, enable a flexible parameterization, resulting in a safety filter that involves a convex quadratic optimization problem, which significantly alleviates the online computational burden. We propose a learning-based approach to synthesize SACBFs. The effect of learning errors on the effectiveness of SACBFs is addressed by constraint tightening and introducing a new concept called contractive-set CBFs. This ensures formal safety guarantees for the learned CBFs and control policies. Simulation results on an inverted pendulum with elastic walls validate the proposed CBFs in terms of constraint satisfaction and CPU time.
@@ I Introduction
Learning-based control methods have demonstrated extensive success across many applications, such as autonomous vehicles [1] and robotics [2]. Both supervised learning and reinforcement learning (RL) have been widely used for controller synthesis. However, learning-based controllers may provide unsafe control actions that result in undesirable or even destructive effects on the system. In control systems, safety means that the trajectories of the closed-loop system should satisfy state and input constraints for the entirety of the system’s evolution. The lack of safety guarantees limits the ability of learning-based control to achieve safety-critical tasks.
There has been an increasing interest in designing controllers for safety-critical systems. Safe RL is a subfield of RL focused on ensuring safety during or after the learning phase. Techniques in safe RL include penalty methods [3], primal-dual policy optimization [4], constraint elimination [5], invariant sets [6], Lyapunov functions [7], and control barrier functions (CBFs) [2, 8]. These methods aim to prevent the agent from taking unsafe actions by either penalizing such actions in the reward function or by modifying the policy to avoid them. However, penalty-based approaches usually faces trade-offs between ensuring safety and achieving optimal performance, while policy refinement to ensure safety typically requires formal safety certificates, which can be challenging to obtain [9].
Model Predictive Control (MPC) is another approach extensively used to ensure safety in control systems [10]. However, traditional online MPC can be computationally intensive for nonlinear constrained systems [11]. Using supervised learning to approximate the MPC policy is an emerging active research topic that can achieve satisfactory performance in both safety and online computational efficiency [12].
Ensuring long-term safety typically requires either long-horizon prediction [13] or the computation of invariant sets [6]. CBFs offer an alternative approach. As CBFs are energy-based functions, their sub-level sets can be used to characterize the long-term safety of a dynamical system, thereby avoiding the complexity associated with invariant-set representations and long-horizon predictions. Given any learning-based controller without safety considerations, an online optimization problem, called a safety filter, is then solved to find the closest safe input to this controller, while safety is gained by adding a constraint based on a CBF into the optimization problem. The safety filter offers a modular approach to impose safety for any pre-existing controller with formal guarantees, and is more online computationally efficient than some other controllers (such as MPC) that need long-term prediction. CBFs provide a theoretically sound way to ensure that long-term safety constraints are respected by a single CBF constraint, while safety filters adjust control inputs in real time to enforce compliance with the CBF constraint. In addition to CBF-based safety filters, there are also predictive safety filters [9] and safety filters that involve projection onto invariant sets [14]. In this paper, we will use the term “safety filter” specifically to refer to a “CBF-based safety filter”.
@@
In the realm of CBF-based methodologies, even though the state and input constraints are known, there is a lack of universally applicable methods for generating valid CBFs, necessitating reliance on manually designed or problem-specific CBFs. For some particular kinds of systems such as linear, piecewise affine [15], or polynomial systems [16], CBFs can be computed by solving convex optimization problems. For nonlinear systems with general constraints, using learning-based algorithms accompanied by advanced function approximators to estimate CBFs has been explored in several contexts. Roughly, there are four kinds of methods to learn CBFs: the optimization-based method [17], the learner-verifier method [18], the Hamilton-Jacobi (HJ) reachability method [19], and the predictive safety filter method [13, 20]. The first two methods can be conservative, i.e., the learned safe set can be a small subset of the maximal controlled invariant set. In comparison, the HJ reachability method relies on the dynamic programming principle to approximate the maximal controlled invariant set iteratively, while the predictive safety filter method is based on a receding-horizon open-loop optimal control problem to implicitly determine the safe set, which also converges to the maximal controlled invariant set as the horizon goes to infinity [21]. A comprehensive comparison of HJ reachability, predictive safety filters, and CBFs can be found in [9].
@@
No matter how the CBF is learned, to control a given system with safety guarantees, it is inevitable to solve an online optimization problem with CBF-based constraints, which is usually non-convex for discrete-time nonlinear systems [22]. To obtain a satisfactory approximation accuracy, sophisticated function approximators such as deep neural networks (NNs) are commonly used to parameterize the CBF [13, 2]. However, this will inherently cause non-convexity and increase the complexity of the optimization problem. As a result, the increased online computational load makes the CBF-based approach unsuitable for situations where fast computation of control inputs is required. Besides, the effect of approximation errors on the validity of the learned CBFs and the resulting control policies has not yet been addressed, as highlighted in [2]. Note that [19] introduces control barrier-value functions by unifying Hamilton-Jacobi (HJ) reachability and CBFs to address issues of finding valid CBFs. However, [19] focuses on theoretical guarantees of robust safety to disturbances but does not address learning errors.
The current paper has the following main contributions:
(i) Computationally efficient safety enforcement using state-action CBFs: We introduce the novel concept of state-action CBFs (SACBFs). Unlike standard CBFs [19, 23, 22], which are often computationally intensive when ensuring safety in nonlinear systems, SACBFs provide a flexible parametrization.The flexibility lies in the ability to enforce safety constraints through a convex quadratic optimization problem, significantly reducing the computational burden during online operations.
(ii) Handling approximation errors in learned CBFs: We propose a constraint tightening approach alongside the novel concept of contractive-set CBFs to address approximation errors in learned CBFs. This ensures that the invariance property of CBFs is maintained even when approximation errors occur, which is not yet achieved in existing work on learning CBFs [19, 17, 13, 24]. Additionally, we examine the relationship between SACBFs and contractive-set CBFs, and develop a new learning-based method to approximate SACBFs. This ensures the safety of the policy filtered by SACBFs, provided the approximation error is sufficiently small. Besides, we discuss the trade-off between online computational efficiency and safety when learning the proposed SACBFs.
@@ II-A Preliminaries
where $x_{t}\in\mathcal{X}\subseteq\mathbb{R}^{n_{x}}$ and $u_{t}\in\mathcal{U}\subseteq\mathbb{R}^{n_{u}}$ are the state and the input at time step $t$, and $f(\cdot,\cdot):\mathcal{X}\times\mathcal{U}\to\mathcal{X}$ is a continuous function satisfying $f(0,0)=0$. We consider a constrained optimal control problem in which the states and inputs should satisfy time-invariant constraints: $x_{t}\in X:=\{x\in\mathcal{X}|h(x)\leq 0\}$ and $u_{t}\in U\subseteq\mathcal{U}$. Here, $h(\cdot):\mathcal{X}\to\mathbb{R}$ is a continuous function that defines the state constraint [^1] For the constraint defined by multiple inequalities $h_{i}(x)\leq 0,\;i=1,2,...,I$, we can let $h(x)=\max_{i\in\{1,...,I\}}h_{i}(x)$. The set $\{x|h(x)\leq 0\}$ is then identical to $\{x|h_{i}(x)\leq 0,\;i=1,2,...,I\}$, and $h$ will be continuous if each $h_{i}$ is continuous.. In most parts of the paper, we assume that $f$ is fully known. However, in Section V.C, we briefly discuss the applicability of our method when there is system uncertainty. Besides, we assume that $X$ is compact and that $U$ is a polytope.
The control objective is to regulate a predefined control policy, which could be a policy learned through various methods such as RL [25], learning-based MPC [14], or any other suitable control technique that needs safety guarantees. The primary concern is ensuring constraint satisfaction after learning, rather than safe exploration [26], which is not the focus here.
@@ II-B Control barrier functions
${\color[rgb]{0,0,1}\forall x\in\mathcal{S}_{B},\;\exists u\in U\text{ s.t. }B(f(x,u))\leq 0.}$ (3)
Furthermore, a CBF is called an exponential CBF if $h(x)\leq B(x),\;\forall x\in\mathcal{S}_{B}$ and if there exists a $\beta\in(0,1)$ such that
${\color[rgb]{0,0,1}\forall x\in\mathcal{S}_{B},\;\exists u\in U\text{ s.t. }B(f(x,u))\leq\beta B(x).}$ (4)
Condition (2) is equivalent to $\mathcal{S}_{B}\subseteq X$, which is usually assumed in literature [22]. With a CBF available, one can generate a safe control policy in $\mathcal{S}_{B}$ by using the following optimization-based approach:
$\pi_{\mathrm{safe}}(x)\!=\!$ $\arg\min_{u\in U}\;||u-\pi_{0}(x)||_{2}$ (5)
@@
$\mathrm{or}\;B(f(x,u))\leq\beta B(x),{\color[rgb]{0,0,1}\mathrm{if}\;B\text{ is an exponential CBF},}$
@@
For general nonlinear systems with state and input constraints, synthesizing a non-conservative CBF is a difficult task. Hand-crafted or application-specific heuristics are mostly used in literature to design candidate CBFs, which can be either unsafe or overly conservative (see [27, Figure 1] for an illustrative example). To deal with this issue, using advanced function approximators to learn a CBF certificate has received much attention (see [2] for a comprehensive survey).
II-C Limitations of existing CBFs
Problem P1: high online computational complexity. One main limitation of (5) is that it needs to solve a usually non-convex optimization problem in real time. If a CBF is represented by a complex function approximator $B_{\theta}(\cdot)$ with the parameter $\theta$, solving (5) may take much online computation time and result in very sub-optimal solutions. Even if a convex $B_{\theta}$ is formed, the constraint in (5) could be non-convex due to the nonlinearity of $f$.
Problem P2: effects of approximation errors. Another limitation of (5) is that the approximation error of $B(\cdot)$ may affect the safety of the system controlled by the optimizer of (5) with $B(\cdot)$ replaced by $B_{\theta}(\cdot)$. Besides, the recursive feasibility of (5) is also not guaranteed with $B_{\theta}(\cdot)$.
@@ III State-action control barrier function
where $Q(\cdot,\cdot):\mathcal{X}\times U\to\mathbb{R}$ is a function of states and actions. We will analyze how to enforce the safety of $\pi_{\mathrm{safe}}$ by imposing conditions on $Q$. To achieve this, we introduce the definition of SACBFs as follows.
@@
A continuous function $Q(\cdot,\cdot):\mathcal{X}\times U\to\mathbb{R}$ is called a state-action control barrier function (SACBF) with a corresponding safe set $\mathcal{S}_{Q}$, if the pair $(Q,\;\mathcal{S}_{Q})$ satisfies
(i) $\mathcal{S}_{Q}$ is non-empty, and $h(x)\leq 0,\;\forall x\in\mathcal{S}_{Q}$,
(ii) $\forall x\in\mathcal{S}_{Q},\;\exists u\in U\text{ s.t. }Q(x,u)\leq 0$, and
(iii) for any $x\in\mathcal{S}_{Q}$, any $u\in U$ that satisfies $Q(x,u)\leq 0$ will make $f(x,u)\in\mathcal{S}_{Q}$.
Unlike the definition of standard CBFs, we do not prescribe the form of $\mathcal{S}_{Q}$ based solely on $Q$. Besides, condition (i) imposes $\mathcal{S}_{Q}\subseteq X$. The following lemma builds the connection between standard CBFs and SACBFs, and provides an explicit form of $\mathcal{S}_{Q}$ when the SACBF is derived from a standard CBF.
@@
(i) For the safety filter (6), any SACBF $Q$ will render $\pi_{\mathrm{safe}}$ safe in $\mathcal{S}_{Q}$. In other words, $\mathcal{S}_{Q}$ is control-invariant.
(ii) If $B$ is a CBF, $Q(\cdot,\cdot)=B(f(\cdot,\cdot))$ will be an SACBF with the safe set $\mathcal{S}_{B}$, i.e., $\mathcal{S}_{Q}=\mathcal{S}_{B}$.
@@
Similarly to designing CBF, the challenge of using (6) is that the explicit form of an SACBF cannot be directly obtained based on Definition 3. We will present a learning-based approach to synthesize SACBFs in Section V.
@@
which will make problem (6) a convex QP with linear and convex quadratic constraints. In (7), $q_{1,\theta}(\cdot):\mathbb{R}^{n_{x}}\to\mathbb{R}$, $q_{2,\theta}(\cdot):\mathbb{R}^{n_{x}}\to\mathbb{R}^{n_{u}}$ and $Q_{3,\theta}(\cdot):\mathbb{R}^{n_{x}}\to\mathbb{R}^{n_{u}\times n_{u}}$ are parameterized functions with all parameters condensed in $\theta$.
If the system (1) has a control-affine form, i.e., $f(x,u)=f_{x}(x)+f_{u}(x)u$, and the linearized system of (1) around the origin is stabilizable, there always exists an SACBF in the form of (7). This is because a quadratic CBF always exists for (1) according to Appendix D, and $Q(\cdot,\cdot)=B(f(\cdot,\cdot))$ is an SACBF if $B$ is a CBF. Additionally, we should note that for general non-affine control systems, a quadratic SACBF is not guaranteed to exist.
Our proposed concept of SACBFs is inspired by the definition of state-action (Q) value functions in RL [25]. Both kinds of functions not only characterize the energy of each state but also quantify the quality of taking each action in each state. In Section V, we will demonstrate that the Q value function for a specific finite-horizon optimal control problem is, in fact, an SACBF.
IV Constraint tightening and contractive-set CBFs
To cope with Problem P2, we propose a method to compute conservative CBFs based on state constraint tightening. In particular, we consider two kinds of conservative CBFs. The first kind includes the CBFs of system (1) under the tightened state constraint $X_{\lambda}:=\{x\in\mathcal{X}|h(x)+\lambda\leq 0\}$ where $\lambda$ is a positive constraint backoff. In this situation, the CBFs still satisfy (3) or (4), while condition (2) is tightened to $h(x)+\lambda\leq B(x),\forall x\in\mathcal{S}_{B}$.
The second one, called the contractive-set CBF, is defined as follows
Definition 4 (Contractive-set CBF)
A CBF $B(\cdot):\mathcal{X}\to\mathbb{R}$ is called a $\lambda$-contractive-set CBF [^2] Note that in fact the definition of contractive-set CBFs given here is not consistent with the common definition of contraction mappings [28]. Formally speaking, the CBF we define should be called a CBF with a contractive safe set [29]. However, for compactness we adopt “contractive-set CBF” in the paper. if there exists a $\lambda>0$ such that